Find the cost per ounce of a sunscreen made from 130 oz of lotion that cost $3.46 per ounce and 90 oz of lotion that cost $11.38 per ounce.
? oz

Answer by Maths68(1474)

130 oz of lotion costs = $3.46 per ounce
90 oz of lotion costs = $11.38 per ounce.
Mixture costs = x per ounce
3.46*130+11.38*90 = (90+130)*x
449.8+1024.2 = 220x
1474=220x
1474/220=220x/220
6.7=x
x=6.7
Mixture costs $6.7 per ounce